Robins Corporation manufactures one product. It does not maintain any beginning or ending Work in Process inventories. The company uses a standard cost system in which inventories are recorded at their standard costs and any variances are closed directly to Cost of Goods Sold. There is no variable manufacturing overhead. The standard cost card for the company’s only product is as follows:

Inputs

Standard Quantity
or Hours

Standard Price or Rate

Standard Cost

Direct materials

3.8

pounds

$

9.50

per pound

$

36.10

Direct labor

0.80

hours

$

18.50

per hour

14.80

Fixed manufacturing overhead

0.80

hours

$

18.00

per hour

14.40

Total standard cost per unit

$

65.30

The standard fixed manufacturing overhead rate was based on budgeted fixed manufacturing overhead of $360,000 and budgeted activity of 20,000 hours.

During the year, the company completed the following transactions:

  1. Purchased 134,700 pounds of raw material at a price of $9.10 per pound.
  2. Used 122,080 pounds of the raw material to produce 32,100 units of work in process.
  3. Assigned direct labor costs to work in process. The direct labor workers (who were paid in cash) worked 26,680 hours at an average cost of $17.20 per hour.
  4. Applied fixed overhead to the 32,100 units in work in process inventory using the predetermined overhead rate multiplied by the number of direct labor-hours allowed. Actual fixed overhead costs for the year were $378,400. Of this total, $297,400 related to items such as insurance, utilities, and indirect labor salaries that were all paid in cash and $81,000 related to depreciation of manufacturing equipment.
  5. Completed and transferred 32,100 units from work in process to finished goods.

Assume that all transactions are recorded on the below worksheet, which is similar to the worksheet shown in your text except that it has been divided into two parts so that it fits on one page. The beginning balances in each of the accounts have been given. PP&E (net) stands for Property, Plant, and Equipment net of depreciation.

1/1

Cash

Raw Materials

Work in Process

Finished Goods

PP&E (net)

=

Materials Price Variance

Materials Quantity Variance

Labor Rate Variance

Labor Efficiency Variance

FOH Budget Variance

FOH Volume Variance

Retained Earnings

a.

$1,000,000

$28,880

$0

$84,890

$566,900

=

$0

$0

$0

$0

$0

$0

$1,680,670

b.

=

c.

=

d.

=

e.

=
Answer #1
Cash Raw Materials Work in Process Finished Goods PP&E (net) = Materials Price Variance Materials Quantity Variance Labor Rate Variance Labor Efficiency Variance FOH Budget Variance FOH Volume Variance Retained Earnings
1/1 1000000 28880 0 84890 566900 = 0 0 0 0 0 0 1680670
a. -1225770 1279650 = 53880
b. -1159760 1158810 = -950
c. -458896 475080 = 34684 -18500
d. -297400 462240 -81000 = -18400 102240
e. -2096130 2096130 =
1833624 = 1833624

Working:

Materials Price Variance = AQ x (AP - SP) = 134700 x ($9.10 - $9.50) = $53880 F
Materials Quantity Variance = SP x (AQ - SQ) = $9.50 x (122080 - 121980) = $950 U
SQ = 32100 units x 3.8 pounds = 121980 pounds
Labor Rate Variance = AH x (AR - SR) = 26680 x ($17.20 - $18.50) = $34684 F
Labor Efficiency Variance = SR x (AH - SH) = $18.50 x (26680 - 25680) = $18500 U
SH = 32100 units x 0.80 hours = 25680 hours
FOH Budget Variance = Budgeted FOH - Actual FOH = $360000 - $378400 = $18400 U
FOH Volume Variance = Budgeted FOH - Applied FOH = $360000 - (25680 x $18) = $360000 - $462240 = $102240 F